A 1.15-kΩ resistor and a 570-mH inductor are connected in series to a 1100-Hz generator with an rms voltage of 14.7 V .

What is the rms current in the circuit?

What capacitance must be inserted in series with the resistor and inductor to reduce the rms current to half the value found in part A?
Answer #1

Voma 14 7 v. R=1.15x1032 L=570810-3H W=25x1100 eads z = R2 + XL2 J (1.15x103)2 + ( W2)2 (115x102)2 + (2x1100x570x103)2 S2 41